# Travel Experience in Public Transport Dataset

### v1

@@ Abstract
The transportation sector holds the potential to change the world towards a greener future if it aligns with increasing mobility needs. One solution is to make public transport an attractive alternative to individual transportation. Real-world data is needed to investigate reasons for and indicators of positive and negative travel experience. Here, we present a GPS-tagged dataset where participants wore an electrocardiogram and reported experience sampling that measured stress, satisfaction, events, and emotions while traveling by tram, train, and bus. An interactive experience map helps to visually explore the data. As benchmark analysis for future users of the dataset, we report significant stress hot spots and satisfaction cold spots during the participants’ journeys. The reported events and emotions, especially in such hot and cold spots, can be analyzed to highlight points of positive and negative travel experience in an ecologically highly valid setting. Data on age and self-identified gender offers insights to differences between user groups. Overall, our dataset enables the combination of qualitative and quantitative methods to identify user’s needs in public transportation.
@@
Transport causes nearly 25% of the world’s carbon dioxide emissions from energy [1]. Buses, railways, and trams constitute 13.8% of kilometers driven by motorized transport on land in the EU in 2020 [2]. The other 86.2% of kilometers were covered by cars [2]. At the same time, public transport transmits more than 50% less greenhouse gases per passenger kilometer than cars when compared to small and medium cars only [3]. One effective step towards worldwide greenhouse gas reduction is, therefore, a shift from private to public transport [4].
@@
While the importance of shifting to public transport is evident, the feasibility of this shift has been the focus of extensive studies. Therefore, a large body of research has investigated factors to enable an increase in public transport usage in terms of efficiency [5, 6], infrastructure [7], and politics [8]. Another important factor for a transition towards greener mobility is the user, who is the ultimate deciding factor in increased public transport usage. Therefore, previous studies have mostly concentrated on time, cost, and reliability aspects that relevantly influence mobility choices ([9, 10, 11]).
Despite the extensive research on technical and practical aspects, the human element - specifically the travel experience and its influence on future travel mode choices - has not been thoroughly explored. Travel experience is closely linked with feelings and emotions [9]. Such feelings and emotions are experienced through a person’s cognitive appraisal and an appropriate physiological response to an event [12]. It is, therefore, essential to use multimodal metrics that consider both the traveler’s emotional state and physiological response [13]. Thus, when assessing travel experiences in intermodal journeys, it is beneficial to integrate subjective assessments with physiological indicators of emotions, as well as relevant contextual data [14, 15]. Other indicators of emotion, like facial expressions or electroencephalography, are not feasible in public transport. While numerous studies have used physiological data to correlate physiological changes with subjective experiences and emotions [16, 17, 18], only first steps have been taken to use physiological data to infer travel experience in public transport [19, 20].
Experience sampling is a method to collect subjective data with high frequency in a ecologically valid setting [21]. By capturing moments of positive and negative experiences as they occur, researchers are provided with data that is less influenced by memory biases in terms of quantity and intensity compared to post-hoc methods [22]. The method has previously been used to assess information need in an automated shuttle [23] and, as a case study, in public transportation of Santiago de Chile [14]. A straightforward method to depict such location-based experience sampling data is through map visualizations. These have previously been used in the automotive sector [24, 25] and for pedestrians and cyclists [26]. Using experience mapping to investigate public transport experience has been suggested by [27] as a method that can pinpoint collective points of negative experience in a transportation system based on the first 23 participants of this dataset. In this paper, we present the complete dataset and include a map visualization.
Overall, the qualitative dimension of the travel experience has been addressed in some studies [28, 29], but synchronous, data-driven datasets are lacking so far. Given this gap, our dataset offers a comprehensive and synchronous approach by collecting data from 44 participants who traveled on a pre-defined route that entailed changes between tram, bus, and train on a 15 km long route. During the journey, we collected global positioning system (GPS) data, heart rate and heart rate variability data, and experience sampling data on satisfaction with travel, stress levels, freely chosen emotion words, and events that occurred. We visualized all these data in an interactive map that enables the researcher to easily investigate the dataset. We identified significant spots of stress and satisfaction and the events and emotions that were reported at these spots as benchmark analysis.
The dataset can be used to develop data analyses methods related to causes for positive and negative travel experience, locations of collective positive or negative travel experiences, and physiological correlates of travel experience during active and passive travel modes. Data on age and self-identified gender can be used to investigate differences between user groups. Overall, our dataset enables the combination of qualitative and quantitative methods to identify methods that measure user’s needs in public transportation.
@@ Methods
  • Getting from home with a freely chosen transportation mode to a central bus station in Brunswick (German city with roughly 250,000 inhabitants),
  • taking a tram to Brunswick main station (GPS location: 52.251332328, 10.537164518),
  • taking a bus (50% of journeys) or a train (the other 50% of journeys) to Wolfenbuettel (German city with roughly 50,000 inhabitants, distance roughly 15km, GPS location: 52.15951774394025, 10.532477626246047),
  • walking to the city center.
and was taken back in reverse. Participants were free to choose to start their journey from Brunswick or Wolfenbuettel. The route was chosen to include as many as possible transportation modes of public transport. Figure 2 displays the route. Data acquisition took place from January to December 2023.
@@
In total, 44 participants (22 self-identified as male, 20 self-identified as female, and two preferred not to fill in any demographic information) took part in the study. They had a mean age of 29.4 years (SD 13.8 years), and none of them indicated to have had any cardiovascular diseases. They indicated to have no specific restrictions during travel except for two persons who indicated to travel with a lot of luggage on all of their journeys. One participant lived in a city with over 500,000 inhabitants, 34 participants lived in a city with up to 500,000 inhabitants, two up to 100,000, four up to 50,000, and one in the countryside. Eighteen had a way to work below 5 km, six had 10-20km, three had 20-30 km, two 30-40km, nine 5-10km, and four less than 1 km. Thirty-four of the participants had a driver’s license. With multiple responses possible, 28 participants usually take a bus to work, 8 drive to work by car, 1 is taken to work by car, 23 cycle to work, 12 take a train to work, 16 a tram and 19 usually walk to work. The participants were acquired from our institute’s participant pool and received 60€-100€ (depending on the number of journeys taken, 10€ per journey) as financial reimbursement for participation. Before starting the study, participants were informed about and agreed to the study’s procedure and the handling of personal data (in line with the General Data Protection Regulation of the EU). The study’s procedure and sharing the data was approved by the institution’s ethical committee (reference no. 07/22).
@@
Subjective measures:
Before the start of the study, participants filled in a demographic questionnaire and the Big Five Inventory (BFI-10) following [30] on a scale from 1 = strongly disagree to 5 = strongly agree with the following questions:
@@
At the beginning of every journey, participants indicated their destination and were asked to indicate every mode change of their journey by pressing the corresponding button in the MovisensXS App (Movisens GmbH in Karlsruhe, Germany, 2023b) on the study smartphone. During their journey, participants were prompted to answer an experience sampling questionnaire every 3.5 minutes using the same app. The decision to adopt a 3.5-minute interval was based on a previous study [20, 32], during which participants were exposed to different frequencies of questionnaire requests. Participants used Likert scales ranging from 1 (indicating low stress and low satisfaction) to 10 (reflecting high stress and high satisfaction) to assess their current satisfaction with the journey and their stress levels. The precise questions asked as translated from German were ’Please rate your current satisfaction with travel’ and ’how stressed do you feel?’. They furthermore answered ’how do you feel independent of your journey?’ from 1 (positive) to 10 (negative), ’How would you describe your current emotion?’, and ’Did anything happen since the last journey?’. If answered with ’yes’, they could choose from the following event categories that were found in a prestudy [20, 32]:
@@
The participants were furthermore free to report a freely chosen emotion word below the choice of categories. We translated the free text inputs to English with Deepl.com. Between the prompted questionnaires, the participants were instructed to freely prompt another questionnaire whenever they encountered something that changed their evaluation or experience of the commute. At the end of every journey, participants rated their overall travel experience using the Satisfaction with Travel Scale (STS) [33]. The STS assesses the emotional and cognitive evaluation of the journey. It comprises of the questions ’How do you feel?’ from Likert scales that measured hurried (1) to relaxed (7), worried (1) to confident (7), stressed (1) to calm (7), tired (1) to alert (7), bored (1) to enthusiastic (7) and annoyed (1) to stimulated (7). Part two of the STS asks ’How do you judge your journey?’ from ’the worst journey I can imagine’ (1) to ’the best I can imagine’ (7), from ’low quality’ (1) to ’high quality (7), and from ’the journey worked out badly’ (1) to ’the trip went smoothly’ (7).
After all journeys, the participants filled out the post-questionnaire containing two decision questions:
  • Will you continue to commute using public transport?
  • If you had to travel from Wolfenbuettel main station to Brunswick main station (or vice versa), would you prefer using the bus or the train?
@@
The participants’ heart rate and heart rate variability (HRV) was once per second extracted from an EcgMove 4 belt from Movisens (Movisens GmbH, Karlsruhe, Germany, 2022).
@@
Throughout their travel, the location was consistently recorded using the movisensXS App and the Movinglab app  [^1] https://movinglab.dlr.de/. While the movisens XS app collected subjective and GPS data, the Movinglab app automatically recognized the travel mode (walk, stationary, train, change, bus, car, bicycle, unknown) and also recorded GPS data. To adhere to the guidelines set by our institute’s data protection committee, participants received explicit instructions to activate the GPS data on the study cell phone solely upon entering the initial major street subsequent to commencing their journey. The GPS signal had a mean accuracy of 8.60m.
Data Preprocessing
The original heart rate data (column HR [1/min]_EcgMove4) was transformed using the median absolute deviation per participant (column ’HR_mad)’ by 1) centering the values by subtracting the participant’s median heart rate from all heart rate data, 2) calculating these values’ median absolute deviation and 3) division of the centered values by the median absolute deviation values (see 0_mad_per_participant.Rmd of the available code). Furthermore, we excluded heart rate values during times that the ECG belt recognized active travel modes like walking or standing, and only included data where participants were sitting for longer than one minute in the column HR_mad_filtered.
@@
The data is available on the Open Science Framework [34]. Each .csv contains the data of one participant. Find each column explained in the tables at the end of his paper.
@@
We include an interactive map  [^2] https://estherbosch.shinyapps.io/affectivemaps that enables dynamic visualization of our dataset with the R Shiny package [35] in R [36]. The data displayed on the map were filtered to only include data where more than two participants contributed to an area. This was done by rasterizing the data into 100mx100m longitude and latitude squares and excluding all squares that contained less than two participants, (see 1_HR_mad_filtered_2squares.Rmd). All relevant variables measured during the journeys can be chosen by a drop-down menu and data of all participants or single participants can be selected, see Figure 3. A click on each data point provides all available information on this point’s stress rating, satisfaction rating, heart rate, and freely reported emotions.
@@
As reference analysis for detecting relevant points of travel experience in public transport with our dataset, we calculated cold spots and hot spots of satisfaction and stress ratings using the Getis-Ord Gi statistic following [37] with 100m radius circles around each data point. This is an analysis to indicate a statistically significant cluster of high or low values in each rating. A statistically significant hot spot exhibits a high value and has neighboring high values. The cumulative value of a feature (the data points within the circle around a data point with a 100m radius, in this case) and its adjacent data points are assessed relative to the cumulative value of all features. When this cumulative value significantly diverges from the anticipated local sum more than expected by random chance, it represents a statistically significant stress or satisfaction rating. To all calculations, we applied Bonferroni correction for multiple tests. The hot and cold spots are displayed on the interactive map as the middle points of the 100m radius circles with statistically significant hot and cold spots. The respective hot- and cold spot reported events within the 100m radius are displayed in Figure 4. There were 28 freely reported emotions in the satisfaction cold spot, of which 11 (39.3%) were positive ("calm", "happy", "joy", "relaxed", "satisfied", "in a good mood and relaxed", "motivated", "excited", "good", "serene", "enjoyable"), also 11 (39.3%) were negative ("upset", "annoyed", "tired and funny to drive past home", "tired", "slightly tense", "tense", "a little annoyed and tired", "annoyed but getting better", "still annoyed", "slightly annoyed", "bored") and 6 (21.4%) neutral ("okay", "neutral", "indifferent", "unchanged", "slow", "quiet"). There were 12 freely reported emotions in the stress hot spot, of which 3 (25%) were positive ("Good","Relaxed", "Satisfied"), 4 (33.33%) were negative ("Stressed", "Tired", "Stressed because the tracked mode does not show up","Annoyed") and 5 (41.67%) neutral ("Quiet", "Okay", "Good, a little stressed because late, a lot to do", "Talkative", "Neutral").
@@
Heart Rate Measures and Permutation Analysis of Heart Rates in Cold-/Hot Spots

### v2

# Travel experience in public transport: Experience sampling and cardiac activity data for spatial analysis

@@ Abstract
The transportation sector has the potential to enable a greener future if aligned with increasing mobility needs. Making public transport an attractive alternative to individual transportation requires real-world data to investigate reasons and indicators of positive and negative travel experiences. These experiences manifest not only in subjective evaluations but also in physiological reactions like cardiac activity. We present a geo-referenced dataset where participants wore electrocardiograms and reported real-time stress, satisfaction, events, and emotions while traveling by tram, train, and bus. An interactive experience map helps to visually explore the data, with benchmark analyses identifying significant stress hot spots and satisfaction cold spots during journeys. Events and emotions in these spots highlight positive and negative travel experiences in an ecologically valid setting. Data on age and self-identified gender provide insights into differences between user groups. Despite including only 44 participants, the dataset offers a valuable foundation for transportation researchers and mobility providers to combine qualitative and quantitative methods for identifying public transportation users’ needs.
@@
Transport accounts for nearly 25% of global carbon dioxide emissions from energy use [1]. In the EU, buses, railways, and trams made up 13.8% of motorized kilometers traveled on land in 2020, while cars dominated with 86.2% of kilometers driven [2]. Public transport, however, produces more than 50% less greenhouse gas emissions per passenger kilometer compared to small and medium-sized cars [3]. Shifting from private to public transport is thus a critical step toward reducing global greenhouse gas emissions [4].
@@
While the need to increase public transport usage is well-researched, the feasibility of this shift has been extensively studied, with research focusing on factors such as efficiency [5, 6], infrastructure [7], and policy interventions [8]. Another important factor in this transition is the user, whose preferences and decisions ultimately drive public transport adoption. Previous studies have primarily addressed practical considerations like time, cost, and reliability that influence mobility choices [9, 10, 11].
However, beyond these practical considerations, the emotional and experiential dimensions of travel remain underexplored. These dimensions are critical, as they influence user satisfaction and long-term loyalty to public transport. Lim et al. [12] introduce the concept of within-trip subjective experiences — cognitive and affective responses to individual travel stimuli mediated by appraisal processes — and propose that these experiences shape travel satisfaction, influence mode-related attitudes and intentions, and potentially modify underlying beliefs about mode choice. Similarly, De Vos et al. [13] demonstrate that attitudes are strongly linked to future travel mode choices. However, existing research on travel experience has largely relied on retrospective questionnaires when inquiring about within-trip subjective experiences. The Satisfaction with Travel Scale (STS) [14] is one such instrument, designed to measure two affective and one cognitive component of travel experience [15]. Olsson et al. [16] were among the first to apply the STS to public transport, analyzing 361 valid questionnaires. Their findings confirmed the factor structure of the STS, with two affective and one cognitive dimension, achieving a Cronbach’s alpha of 0.78. Lunke [17] extended this work by analyzing responses from 14,015 participants using a three-item questionnaire derived from the STS. This study revealed that short waiting times and reliable transport are key drivers of travel satisfaction.
In addition to the STS, alternative methods have been employed to investigate travel experience. Beck et al. [18] developed a best-worst scaling questionnaire to assess attitudes toward aspects of bus trips, including noise levels and the availability of information. Soza Parra et al. [19] implemented an in-situ measurement approach, assessing satisfaction per travel leg on a seven-point scale at four metro stations in Santiago de Chile. This real-time approach reduces reliance on retrospective recollection. Meanwhile, Van Lierop [21] conducted a literature review to identify factors influencing public transport loyalty, highlighting attributes such as cleanliness, comfort, operator helpfulness, and schedule frequency. Ingvardson et al. [22] analyzed 44,956 questionnaires from six European cities using structural equation modeling, identifying accessibility, comfort, staff behavior, and safety/security as the top four dimensions affecting public transport satisfaction.
Despite these advancements, the research remains limited in scope, particularly regarding real-time, multimodal investigations of travel experience. Travel experiences are inherently tied to emotions, which involve both cognitive appraisals and physiological responses [23]. Retrospective methods often fail to capture the dynamic and situational aspects of these experiences, as they are subject to memory biases that can distort both the intensity and frequency of reported events [24]. Experience sampling offers a promising alternative, enabling the collection of subjective data with high temporal resolution in ecologically valid settings [25]. This method minimizes memory bias by capturing moments of positive and negative experiences as they occur, providing richer and more accurate insights compared to post-hoc approaches [26].
Physiological data, such as cardiac activity as measured by heart rate and variability, can complement subjective reports by providing objective indicators of emotional states. Previous studies have successfully correlated cardiac activity changes with subjective experiences and emotions [27, 28, 29, 30, 31]. Especially Heart Rate Variability (HRV) has been found to correlate with sympathetic nervous system activity [32, 33] and therefore higher arousal emotions [34], cortical arousal [35], and stress [36]. Castro et al. [37] incorporated cardiac activity measurements into a model of travel experience, presenting a case study with one participant to demonstrate the feasibility of such methods. Additionally, location-based visualizations of experience sampling data, such as maps, have been used to analyze collective travel experiences. These methods have been applied in contexts like automotive travel [38, 39] and pedestrian or cyclist experiences [40]. Bosch et al. [41] proposed using experience mapping to investigate public transport experiences, suggesting its potential to pinpoint collective pain points in transportation systems.
Despite the recognized importance of emotions and experiences in shaping public transport satisfaction, existing research relies heavily on retrospective and subjective data collection methods. These methods are prone to memory biases and fail to capture the dynamic, real-time aspects of travel experiences. Furthermore, the integration of subjective, cardiac activity, and contextual data in public transport research remains underexplored. This paper addresses these gaps by presenting a comprehensive dataset collected from 44 participants who traveled on a predefined 15 km intermodal route involving tram, bus, and train transitions. The dataset includes global positioning system (GPS) data, heart rate and heart rate variability measurements, and real-time experience sampling data on travel satisfaction, stress levels, self-reported emotions, and significant events. The data is visualized in an interactive map, enabling the identification of collective hot spots of stress and satisfaction and the associated events and emotions. By integrating subjective, cardiac activity, and contextual dimensions, this study provides a synchronous and holistic approach to understanding travel experiences.
This dataset facilitates the development of methods for analyzing the causes of positive and negative travel experiences, identifying collective pain points, and investigating cardiac activity correlates of experiences during intermodal transportation. Additionally, demographic data on age and self-identified gender supports analyses of group-specific differences, enabling a refined understanding of user needs in public transportation.
@@ Methods
  • Getting from home with a freely chosen transportation mode to a central bus station in Braunschweig (German city with roughly 250,000 inhabitants),
  • taking a tram to Braunschweig main station (GPS location: 52.251332328, 10.537164518),
  • taking a bus (50% of journeys) or a train (the other 50% of journeys) to Wolfenbüttel (German city with roughly 50,000 inhabitants, distance roughly 15km, GPS location: 52.15951774394025, 10.532477626246047),
  • walking to the city center (Schlossplatz Wolfenbüttel, GPS location: 52.163224559025245, 10.531477397957316).
and was taken back in reverse. Participants were free to choose to start their journey from Braunschweig or Wolfenbüttel. The route was chosen to include as many as possible transportation modes of public transport. Figure 2 displays the route. Data acquisition took place from January to December 2023.
Figures $bs_{w}f_{r}oute.png,Study_{P}rocedure.png$ go here
@@
In total, 44 participants over the age of 18 years (23 self-identified as male, 21 self-identified as female) took part in the study. The participants were recruited from our institute’s participant pool, which includes individuals who have previously signed up to participate in studies. To mitigate potential recruitment biases, we ensured that the sample was balanced by gender. Participants were selected based on their availability and willingness to participate, and no specific demographic quotas beyond gender balance were applied. Their commute distances, size of the city they live in, education levels and commute transport modes, educational background are shown in Tables 1, 2, 3, and 4. Their answers to the Big Five Inventory (BFI-10), and to the Affinity to Technology (ATI-9) scale are provided in the columns ’BFI’ and ’ATI’. They had a mean age of 29 years (SD 14 years), and none of them indicated to have had any cardiovascular diseases. They indicated to have no specific restrictions during travel except for two persons who indicated to travel with a lot of luggage on all of their journeys. The participants received 60€-100€ (depending on the number of journeys taken, 10€ per journey) as financial reimbursement for participation. Participants who did not hold a yearly regional public transport ticket were compensated for their travel expenses during the study. Before starting the study, participants were informed about and agreed to the study’s procedure and the handling of personal data (in line with the General Data Protection Regulation of the EU, approved by the institution’s data protection and legal committee). The study’s procedure was approved by the institution’s ethical committee (reference no. 07/22).
Tables 1-4 go here
@@
In this study, subjective travel experience was assessed across three distinct time frames: (1) once per participant (e.g., demographic data, personality traits measured by the Big Five Inventory, and technology affinity via the Affinity for Technology Interaction scale); (2) once per journey (e.g., destination details and an overall satisfaction evaluation); and (3) during travel, using experience sampling questionnaires administered at 3.5-minute intervals. This multi-method approach integrates real-time data with post-travel assessments and mode choice questionnaires, enabling an investigation of decision-making heuristics [24] and their relationship with during-travel experiences, post-journey satisfaction with travel (STS), and future mode choice behaviors.
This methodology combines the strengths of experience sampling and retrospective questionnaires. Experience sampling enhances ecological validity, captures within-person variability, and reduces memory bias [26]. On the downside, the highly regular time intervals of reporting experience will make participants more aware of their internal states, which in turn might change them [42]. In contrast, retrospective questionnaires avoid repeated assessments and therefore minimize atypical focus on internal states and provide a broader contextual overview [26]. Furthermore, the inclusion of physiological data adds an objective dimension to the subjective real-time experience, capturing multimodal aspects of travel experiences, as proposed by Scherer’s dynamic framework [23]. These physiological measures complement self-report data, which can be constrained by social desirability, cognitive biases, cultural norms, and the limitations of numerical scales [26].
Contextual factors like time of day and year, weather and delays significantly influence travel satisfaction, shaping both real-time experiences and post-travel attitudes. With the provided GPS and timestamp data, it is possible to obtain weather information and delays. Drawing on Kaiser’s framework [43], which emphasizes the interaction of attitudes, behavior intentions, and external constraints, these situational factors mediate the relationship between during-travel experiences, satisfaction, and future transport choices. Physiological stress markers (e.g., heart rate) reveal how these stressors manifest physically, while subjective self-reports capture individual perceptions influenced by personal attitudes. Integrating real-time contextual data, such as the weather, with subjective and physiological measures provides a comprehensive understanding of how external constraints impact travel experiences.
Subjective measures - once per participant:
Before the start of the study, participants filled in a demographic questionnaire indicating their usual commute distance and mode, the size of the city they live in and education level. They also answered the Big Five Inventory (BFI-10) following [44] on a scale from 1 = strongly disagree to 5 = strongly agree with the following questions:
@@
At the end of all their journeys, participants answered the following questions with free text input:
  • What was the main stressor for you during commuting?
  • Which events influenced you during commuting ? Name up to five positive and up to five negative events.
  • Will you continue to commute using public transport? If no: What would have to get better for you to use public transport for commuting? If yes: why?
  • If you had to travel from Wolfenbüttel main station to Braunschweig main station (or vice versa), would you prefer using the bus or the train? Explain your answer.
Subjective measures - once per journey:
At the beginning of every journey, participants indicated their destination and were asked to indicate every mode change of their journey by pressing the corresponding button in the MovisensXS App (Movisens GmbH in Karlsruhe, Germany, 2023b) on the study smartphone. At the end of every journey, participants rated their overall travel experience using the Satisfaction with Travel Scale (STS) [14]. The STS assesses the emotional and cognitive evaluation of the journey. It comprises of the questions ’How do you feel?’ from Likert scales that measured hurried (1) to relaxed (7), worried (1) to confident (7), stressed (1) to calm (7), tired (1) to alert (7), bored (1) to enthusiastic (7) and annoyed (1) to stimulated (7). Part two of the STS asks ’How do you judge your journey?’ from ’the worst journey I can imagine’ (1) to ’the best I can imagine’ (7), from ’low quality’ (1) to ’high quality (7), and from ’the journey worked out badly’ (1) to ’the trip went smoothly’ (7).
Subjective measures - during travel:
During their journey, participants were prompted to answer an experience sampling questionnaire every 3.5 minutes using the same app. The decision to adopt a 3.5-minute interval was based on a previous study [46, 47], during which participants were exposed to different frequencies of questionnaire requests. Participants used Likert scales ranging from 1 (indicating low stress and low satisfaction) to 10 (reflecting high stress and high satisfaction) to assess their current satisfaction with the journey and their stress levels. The precise questions asked as translated from German were ’Please rate your current satisfaction with travel’ and ’how stressed do you feel?’. They furthermore answered ’how do you feel independent of your journey?’ from 1 (positive) to 10 (negative), ’How would you describe your current emotion?’ (free text input), and ’Did anything happen since the last journey?’. If answered with ’yes’, they could choose from the following event categories that were found in a prestudy [46, 47]:
@@
If they chose ’report a different category’, a free text input opened up. We translated the free text inputs of the ’emotion’ and the ’event’ free text input to English with https://www.deepl.com/de/translator. Between the prompted questionnaires, the participants were instructed to freely prompt another questionnaire whenever they encountered something that changed their evaluation or experience of the commute.
@@
The participants’ electrocardiogram was recorded once per second and extracted from an EcgMove 4 belt from Movisens (Movisens GmbH, Karlsruhe, Germany, 2022). The Movisens Data Analyzer Version 1.13.27 was used to extract heart rate and different measures of heart rate variability in time and frequency domain. The data that is read into the R-Markdown code ’01_Basics-read-translate.Rmd’ stems from this Data Analyzer set to an output rate of 1Hz. Leys et al. [48] suggest that the median absolute deviation (MAD) is a more robust measure of variability than commonly used z-scores. Therefore, the original heart rate data (column ’ecgbelt_hr’) was transformed using the median absolute deviation per participant (column ’ecgbelt_hr_mad’) by 1) centering the values by subtracting the participant’s median heart rate from all heart rate data, 2) calculating these values’ median absolute deviation and 3) division of the centered values by the median absolute deviation values (see 02_Preprocessing.Rmd of the available code). We excluded MAD heart rate values above 3 and below -3 following [49, 48]. The same MAD-correction was done for heart rate variability as measured by root mean squared differences and is saved in the column ’ecgbelt_hrv_mad’.
Separating changes in cardiac activity due to physical activity and emotion is a difficult task [30]. Previous research indicates that the initial phase of heart rate recovery, particularly following low-intensity activities like slow walking, occurs within the first minute after transitioning to a seated position. This rapid adjustment reflects the activation of the parasympathetic nervous system and the reduction in sympathetic activity, facilitating a return to near-resting heart rate levels [50]. Therefore, besides the complete heart rate and HRV data column, we provide a column where we excluded heart rate values during times that the ECG belt recognized active travel modes like walking or standing, that include data where participants were sitting for longer than one minute (see column ’ecgbelt_hr_mad_filtered’).
@@
Throughout their travel, the time and location was consistently recorded using the movisensXS App and the Movinglab App (https://movinglab.dlr.de/). While the movisens XS app collected subjective and GPS data, the Movinglab App automatically recognized the travel mode (walk, stationary, train, change, bus, car, bicycle, unknown). To adhere to the guidelines set by our institute’s data protection committee, participants received explicit instructions to activate the GPS data on the study cell phone solely upon entering the initial major street subsequent to commencing their journey. The GPS signal had a mean accuracy of $8\pm 12m$.
@@
The data is available on the Open Science Framework at https://osf.io/rgkvq under a CC-By Attribution 4.0 International license. Each participant’s data file (.csv) is saved in the folder ’data.zip’. ExperienceMapsDataset.csv contains the complete dataset; ExperienceMapDataset_over2 includes only areas where more than two participants contributed within a 100x100m region; and ExperienceMapDataset_over2_10sec provides a downsampled version at 0.1 Hz. As reading in the large data files often omits data from the final questionnaire only containing data at the very end, we also uploaded the original .xlsx-data files of the experience sampling that is not yet merged with the GPS and physiological data by the Movisens Data Analyzer in the folder ESM_xlsx.zip. Find each column explained in Tables 5-11 at the end of his paper.
@@
We collected 5144 ESM questionnaire answers with an average of $117\pm 72$ answers per participant. This corresponds to an average response rate of $86.82\pm 13.86\%$. The number of answered ESMs varied from 18 to 501 and the response rate from 50.2% to 100%. Participants traveled in total $547\pm 209$ minutes over all trips as indicated by their experience sampling. The participant who traveled the shortest only traveled 261 minutes in total, the longest 1498 minutes (most likely due to a much too late press of the ’end travel’ - button). Participants took at least five and at most 12 trips in total with a mean of $9\pm 2$ trips. One trip took on average $59\pm 26$ minutes over all participants. Figure 3 displays the overall reported events’ frequency. Figure 5 shows that ratings did not vary significantly across the different hours of the day or months of the year. Only November stands out with generally lower satisfaction levels and higher stress levels. Notably, November 2023 experienced 210% of the usual precipitation for that month in Braunschweig as indicated at https://www.wetterkontor.de/de/wetter/deutschland/monatswerte-station.asp?id=10348&yr=2023&mo=-1. In contrast, December, which had an even higher precipitation level at 287% of the monthly average, displayed higher satisfaction levels. This discrepancy may be attributed to the nature of the precipitation: with an average temperature of 5°C in December, it is likely that much of the precipitation fell as snow rather than rain, as in November.
Figure $seasons.pdf$ goes here Figure $EventcatsColdhotspots.pdf$ goes here
The accuracy of the GPS data was $8\pm 12$ meters overall as indicated by the GPS measurement of the Movisens XS App.
The travel mode as reported by the participants themselves includes 101.32 hours of data collected in a bus, 48.82 hours collected in the train, 41.83 hours collected in the tram, 76.90 hours of waiting time and 50.86 hours of walking.
Self-reported travel modes agrees with the DLR Keepmoving App with a Cohen’s Kappa of .711 (88% agreement) of data points, and the Movisens ECG Move chest belt agrees with a Cohen’s Kappa of .11 ( 19% agreement) with the self-reported travel modes. Figure 6 displays one exemplary mode plot for journey 1 of participant 1 for self-reported mode, mode detection by the Movisens chest belt and mode detection by the DLR App movisens. Figure 8 displays the confusion matrix between the travel mode recognition types for the mode types that they overlapped.
Figures moderec.pdf and conf.pdf go here.
[…]
Figures $HR_{m}ode.pdf$ and activeHR.pdf go here.
The satisfaction with the journey by event is shown in Figure 4. On average, the begative events ’negative driving behavior’, ’missing infrastructure’, ’disturbing people’, ’physical discomfort’ and ’delay’ have an average travel satisfaction rating of $6.48\pm 2.08$. In the positive events ’connection reached’, ’positive interaction’, ’media entertainment’, ’physical comfort’ and ’beautiful surroundings’ have an average satisfaction rating of $8.14\pm 1.68$.
Figure eventsatis.pdf goes here.
We include an interactive map at https://estherbosch.shinyapps.io/affectivemaps that enables dynamic visualization of our dataset with the R Shiny package [51] in R [52]. The data displayed on the map were filtered to only include data where more than two participants contributed to an area in order to deceive the participant’s home street in case they started the GPS tracking earlier than instructed. This was done by rasterizing the data into 100mx100m longitude and latitude squares and excluding all squares that contained less than two participants. All relevant variables measured during the journeys can be chosen by a drop-down menu and data of all participants or single participants can be selected, see Figure 10. A click on each data point provides all available information on this point’s stress rating, satisfaction rating, heart rate, and freely reported emotions.
Figure shiny.PNG goes here.
@@
We calculated cold spots and hot spots of satisfaction and stress ratings using the Getis-Ord Gi statistic following Getis [53] with 100m radius circles around each data point. This is an analysis to indicate a statistically significant cluster of high or low values in each rating. A statistically significant hot spot exhibits a high value and has neighboring high values. The cumulative value of a feature (the data points within the circle around a data point with a 100m radius, in this case) and its adjacent data points are assessed relative to the cumulative value of all features. When this cumulative value significantly diverges from the anticipated local sum more than expected by random chance, it represents a statistically significant stress or satisfaction rating. To all calculations, we applied Bonferroni correction for multiple tests. The resulting hot spots and cold spots are visualized on an interactive map, showing all data points within the 100-meter radius and highlighting statistically significant areas. However, it is important to interpret these results cautiously given the limited number of data points. With the small sample size, the results may be overfitted or less generalizable, and the identified hot and cold spots could be influenced by outliers or limited regional variation. Data from 34 out of 44 participants contributed to the satisfaction cold spot analysis and fourteen out of 44 participants reported their stress levels within the identified stress hot spots, contributing to the calculation of these locations.
One satisfaction cold spot was located at the main interchange station in Braunschweig, where participants were on their way from or to home. This station is often crowded, and finding the correct tram station can be challenging due to the multiple options available. The second satisfaction cold spot was found in the center of Wolfenbüttel, approximately 500 meters east of the designated start and end point of the journey. At the end of the study several participants reported frustration about having to travel all the way to Wolfenbüttel center only to immediately turn back and begin their return journey. While the 500-meter offset is unexplained, this area therefore roughly aligns with expectations as a satisfaction cold spot. No statistically significant satisfaction hot spots were identified. The reported events within the cold spot are displayed in Figure 3. 28 freely reported emotions in the satisfaction cold spots, of which 11 (39.3%) were positive as evaluated by the first author ("calm", "happy", "joy", "relaxed", "satisfied", "in a good mood and relaxed", "motivated", "excited", "good", "serene", "enjoyable"), also 11 (39.3%) were negative ("upset", "annoyed", "tired and funny to drive past home", "tired", "slightly tense", "tense", "a little annoyed and tired", "annoyed but getting better", "still annoyed", "slightly annoyed", "bored") and 6 (21.4%) neutral ("okay", "neutral", "indifferent", "unchanged", "slow", "quiet"). The mean satisfaction in the satisfaction cold spots is $6.71\pm 1.60$, compared to $7.67\pm 1.90$ in all data.
One of the stress hot spots is located along a segment of the route that falls outside the official study route, making its origin unclear. This hot spot corresponds to a street serviced only by buses, and it is likely linked to several participants’ journeys, either heading to or returning from their homes. The second stress hot spot is located on the train route just before Wolfenbüttel. At this location, all participants were traveling by train. It is possible that stress levels were influenced by concerns about catching the train back while still having to walk to Wolfenbüttel city center. No stress cold spots were identified. The reported events within the hot spot are displayed in Figure 3. There were 12 freely reported emotions in the stress hot spots, of which 3 (25%) were positive ("Good","Relaxed", "Satisfied"), 4 (33.33%) were negative ("Stressed", "Tired", "Stressed because the tracked mode does not show up","Annoyed") and 5 (41.67%) neutral ("Quiet", "Okay", "Good, a little stressed because late, a lot to do", "Talkative", "Neutral"). The mean stress rating in the stress hot spots was $5.32\pm 2.24$ compared to a mean stress rating of $2.70\pm 1.70$ overall.
@@
0.0.1 Heart Rate Measures and Permutation Analysis of Heart Rates in Cold and Hot Spots